Please answer the following questions that a person new to this course would be able to understand.( Include theorem.)

Problem 6: Consider the linear systems of differential equations a) Sketch the direction Seld for the line gystem. write StreamPlotl(x-2y, 2x-3y] İn Wolframı Alpha a) Use the method of elimination to find a second order linear differential equation that is satisfied by (t b) Find particular solutions x(t) and y(t) such that x(0) 1 and y(0) 2
Answer #1

Given the linear systems of differential equations
   2 3
The sketch of the direction field (v_x, v_y) in this case is
  

a)

   We use the method of elimination by taking one more time derivative of the 1st equation of x(t)
     x''= x'-2y'
Now in order to eliminate y'(t), we use the 2nd equation
   x''= x'-2(2x-3y)
And so,
  x''= x'-4x+6y
And from the first equation, we replace y, i.e.,
   x'=x-2y\Rightarrow 2y=x-x'
And this implies
x''= x'-4x+3(x-x')
  \Rightarrow x''= -2x'-x
   \Rightarrow x''+2x'+x=0 ................(1)
This is the second order differential equation.

b)

  And let us assume that the ansatz for the solution is
   x(t)=Ae^{\alpha t}
And so, putting this in the equation (1), we get
  A(\alpha^2+2\alpha +1)e^{\alpha t}=0
  \Rightarrow \alpha^2+2\alpha +1=0
   \Rightarrow \alpha=-1,~-1
So, the two roots are equal. And so, the general solution is
   x(t)=(A+Bt)e^{\alpha t}=(A+Bt)e^{-t}
where, A and B are two undetermined constants which are to be fixed by the initial conditions.

  The initial conditions are
   x(0)=1
And
   x'(0)=x(0)-2y(0)=1-(2\times 2)=-3
  \Rightarrow x'(0)=-3

The first condition x(0)=1 implies
  x(0)=A\Rightarrow A=x(0)=1
And taking the derivative of the solution x(t), we get
x'(t)=-(A+Bt)e^{- t}+Be^{-t}
  \Rightarrow x'(t)=\left (B-(A+Bt) \right )e^{-t}
And so, the 2nd initial condition implies
   x'(0)=\left (B-A \right )
Now given
  \Rightarrow x'(0)=-3
So, we get
   B-A =-3
Nw we have already got, A = 1, so, we get
   B =A-3=1-3
  \Rightarrow B =-2
And so, the solution is
x(t)=(1-2t)e^{-t}
And so,
   x'(t)=-(1-2t)e^{-t}-2e^{-t}
  \Rightarrow x'(t)=-(3-2t)e^{-t}
And so, from the first equation,
   x'(t)=x(t)-2y(t)
  \Rightarrow y(t)=\frac{1}{2}\left (x(t)-x'(t) \right )
So, as we have already got an answer for x(t) and x'(t), so, we get
   \Rightarrow y(t)=\frac{1}{2}\left ((1-2t)e^{-t}+(3-2t)e^{-t}\right )
  \Rightarrow y(t)=\frac{1}{2}\left (1-2t+3-2t\right )e^{-t}
  \Rightarrow y(t)=\frac{1}{2}\left (4-4t\right )e^{-t}
  \Rightarrow y(t)=2\left (1-t\right )e^{-t}
This is the solution for y(t).